### v1

@@ Abstract
Perspective distortion (PD) causes unprecedented changes in shape, size, orientation, angles, and other spatial relationships of visual concepts in images. Precisely estimating camera intrinsic and extrinsic parameters is a challenging task that prevents synthesizing perspective distortion. Non-availability of dedicated training data poses a critical barrier to developing robust computer vision methods. Additionally, distortion correction methods make other computer vision tasks a multi-step approach and lack performance. In this work, we propose mitigating perspective distortion (MPD) by employing a fine-grained parameter control on a specific family of Möbius transform to model real-world distortion without estimating camera intrinsic and extrinsic parameters and without the need for actual distorted data. Also, we present a dedicated perspectively distorted benchmark dataset, ImageNet-PD, to benchmark the robustness of deep learning models against this new dataset. The proposed method outperforms on existing benchmarks, ImageNet-E and ImageNet-X. Additionally, it significantly improves performance on ImageNet-PD while consistently performing on standard data distribution. Further, our method shows improved performance on three PD-affected real-world applications: crowd counting, fisheye image recognition, and person re-identification. We will release source code, dataset, and models for foster further research.
@@
Perspective Distortion Self-supervised Learning Robustness
@@ 1 Introduction
Perspective distortion (PD) in real-world imagery is omnipresent and poses challenges in the development of computer vision applications. PD arises from various factors, including camera positioning, depth, intrinsic parameters such as focal length and lens distortion, and extrinsic parameters such as rotation and translation. These factors collectively influence the projection of 3D scenes onto 2D planes [34], altering semantic interpretation and local geometry. Precisely estimating these parameters for perspective distortion correction is a challenging task, presenting a critical barrier to developing robust computer vision (CV) methods.
@@
Previous studies in distortion correction predominantly concentrated on the correction of PD, with limited emphasis on enhancing the robustness of CV applications. Consequently, these correction methods transform CV tasks, such as image recognition and scene understanding, into a two-stage process: firstly, rectifying the distortion, and secondly, engaging in task-specific learning. GeoNet[23] employs CNN to predict distortion flow in images without prior distortion-type knowledge. PCL[46] uses perspective crop layers in CNNs to correct perspective distortions for 3D pose estimation. A cascaded deep structure network[37] corrects wide-angle portrait distortions without calibrated camera parameters. Another study[49] predicts per-pixel displacement for face portrait undistortion. PerspectiveNet[20] and ParamNet[20] predict perspective fields and derive camera calibration parameters, respectively. Methods for fisheye image distortion correction are also developed[43, 45].
@@
Since the correction methods are fundamentally not focused on robust representation learning for CV tasks, they essentially follow an inefficient two-step process. One possible approach is to focus on developing task-specific methods where PD is noticeable, such as crowd counting[44, 17], object re-identification[2], autonomous vehicle[41], and recognizing texts in scene[31]. These task-specific approaches fall short in providing generalizable and robust representation learning for PD mitigation, a crucial step for advancing CV research towards distortion robustness.
@@
The core reason behind the lack of generalizable and robust representation learning CV methods to mitigate PD is due to the absence of a real-world PD-affected training dataset. Both capturing real-world PD data and synthesizing PD-affected image data using parameter estimation are far from reality. Although existing benchmarks (ImageNet-X [18], ImageNet-E [22]) are partially affected by perspective distortion, a dedicated benchmark is needed. Therefore, there is a need to artificially synthesize PD in existing data without estimating camera intrinsic and extrinsic parameters. Since PD is non-linear in nature, the focus should be shifted from affine to potential non-linear transformation.
@@
One potential non-linear family of transformation is Möbius transform [1, 30] in complex theory. Due to its conformal nature (preserving angles during transformation), [50] utilized the Möbius transform for eight specific transformations, including various twists and spreads, as a data augmentation strategy to enhance model performance on standard data distributions. In other works, Möbius convolution[29], the authors design a Möbius-equivalent convolution operator in spherical CNN for shape classification and image segmentation tasks. OmniZoomer[5] integrates Möbius transformations into a neural network, enabling movement and variable zoom on omnidirectional images while addressing the challenges of blur and aliasing associated with such transformations. MöbiusGCN[3] addresses the limitation of existing graph convolution network GCNs in explicitly encoding transformations between joints in 3D human pose estimation. None of these works focus on PD.
@@
Due to the non-linear characteristics of Möbius transforming geometry in complex space while ensuring conformality, it is a potential approach to imitate perspective distortion when adapted to image data. Therefore, in the proposed mitigating perspective distortion (MPD) method, we mathematically model PD in the images by employing fine-grained parameter control on Möbius. It is governed by four parameters $a,b,c$ and $d$ (Eq. 1). MPD focuses on a specific family of transforms corresponding to the parameter $c$ to artificially synthesize PD. The core contribution of our method, MPD, is controlled parametrization for orientation (Fig. 1) and varying intensity (Fig. 2), which is capable of representing real-world perspective distortion.
@@
MPD transforms input image coordinates into complex vectors and performs a PD-specific family of Möbius transforms. Transformed complex vectors are then remapped to real-valued pixels, later discretized and interpolated to obtain images mimicking perspective distortion while maintaining the integrity of visual concepts. Notably, MPD achieves this without estimating intrinsic and extrinsic parameters, thus bypassing the need for actual distorted data.
@@
Incorporating PD in deep learning is beneficial for robustness, since it can be used as an augmentation resulting in improved performance on different computer vision tasks. To address the non-availability of a suitable benchmark, we have developed a dedicated perspectively distorted benchmark dataset, ImageNet-PD, derived from the ImageNet[9] validation set to evaluate the robustness of deep learning models against perspective distortion. We showcase a lack of robustness of standard deep learning models against perspective distortion by benchmarking them on ImageNet-PD.
@@
We employ ImageNet-PD in our proposed MPD method and extensively investigate it across supervised and self-supervised learning methods (SimCLR [7], DINO [6]). MPD demonstrates improved performance on existing benchmarks, ImagNet-E [22], and ImageNet-X [18]. MPD improves $10\%$ on ImageNet-PD while consistently maintaining the performance on the original ImageNet validation set. MPD shows performance improvements when adapted to multiple real-world CV applications. Specifically, MPD advances crowd counting with a novel method, MPD-AutoCrowd, achieving the highest performance (to the best of our knowledge) with a mean absolute error of 50.81 on ShanghaiTech-Part-A[47] and 96.80 on UCFCC50[16]. It demonstrates effective knowledge transfer in fish-eye image classification, improving performance by $3\%$ on the VOC-360[11] dataset. Also, MPD, combined with a transformer-based self-supervised method, Clip-ReIdent[12], excels in person re-identification, achieving mean average precision of 97.02 and 98.30 with and without re-ranking, respectively, on the DeepSportradar dataset[38].
The main contributions of this paper are as follows:
@@
    We develop a dedicated perspectively distorted dataset (ImageNet-PD) derived from the ImageNet[9] validation set to benchmark the robustness of deep learning models against PD to address the non-availability of a suitable benchmark.
@@ 2 Methodology
Conformality: Conformal transformations preserve angles between curves. The angle between two curves at a point is given by the argument of the complex ratio[14] of their derivatives. Let $f(z)$ and $g(z)$ represent curves in complex space. Their Möbius transforms are $\Phi(f(z))$, and $\Phi(g(z))$. If transformation preserves this ratio, it preserves angles. Eq. 3 shows that the angles are preserved, hence conformal [30]. Details are in suppl. material (5).
@@ 2.1 Mitigating perspective distortion (MPD)
The proposed MPD method mathematically models perspective distortion in the image using a family of Möbius transform. First, it maps the discrete coordinates $(x,y)$ of the input image $I$ from pixel space to complex vector $z=x+iy$ in complex space. Second, it applies Möbius transform to $z$ and achieves corresponding non-linear transformed representation $z_{m}=x_{m}+iy_{m}$ by notably controlling real and imaginary parts of parameter $c$ in Eq. 1. Next, the transformed complex vector $z_{m}$ is mapped to real-valued pixel space $(x_{m},y_{m})$, which is discretized to get discrete pixel coordinates $(x_{d},y_{d})$. Obtained discrete pixel coordinates are interpolated to achieve smoother output transformed images $I_{MPD}$ mimicking perspective distortion. In this manner, MPD can synthesize perspective distortion artificially without estimating intrinsic and extrinsic parameters of perspective distortion. Noticeably, It eliminates the need for perspectively distorted real image data.
@@
The core part of MPD in modeling the PD lies in controlling the transformation through parameter $c$. Parameter $c$, being a complex number has real ($c_{real}$) and imaginary ($c_{imag}$) components; it is noteworthy that these components introduce real-world varieties of PD in the input image. $c_{real}$ perspectively distorts the in input image horizontally whereas $c_{imag}$ perspectively distorts the in input image vertically. The sign of $c_{real}$ dictates the direction ($+ve$ for left view and $-ve$ for right view) of distortion, whereas magnitude defines the intensity of applied distortion. Similarly, the sign of $c_{imag}$ dictates the direction ($+ve$ for top view and $-ve$ for bottom view) of distortion, whereas magnitude defines the intensity of applied distortion. The impact of signs of $c_{real}$ and imaginary $c_{imag}$ is depicted in Fig. 1. The synthesis of the increasingly distorted left view of the input image is demonstrated in Fig. 2 by controlling the intensity of $c_{real}$ component of MPD. Demonstrations of increasingly distorted right, top, bottom, and other views are provided in Fig. 15 in the suppl. material. The main contribution of our MPD method is its ability to control parameterization for varying intensity and orientation, effectively representing real-world perspective distortions.
One of the ways to incorporate MPD is augmentation, which is investigated in this work. MPD is effortlessly employable to incorporate into the learning process to make networks robust against perspective distortion, materializing the proposed approach to shifting the focus from exclusion to inclusion of PD. The above-described method MPD is presented in Algorithm 1. MPD returned a transformed image of a ’cat’ as shown in step 11. Four regions are marked with colors (green, red, blue, and yellow) to observe the correspondence between input and transformed image.
Although the current work does not focus on other parameters ($a$, $b$, and $d$), their impact on transformation is worth mentioning. The real component of parameters $a$ and $d$ controls scaling, and the imaginary components control rotations. The real component of parameter $b$ controls horizontal translations, and the imaginary component controls vertical translations. Their visualizations are provided in Fig. 16 in the suppl. material. The transformation parameters are set as: $a=1.0+0.0i$, $b=0.0+0.0i$, $d=1.0+0.0i$. The real and imaginary components of $c$ range from $0.2$ to $0.3$ to mimic perspectiveness. These values are uniformly applied in all experiments unless otherwise stated. The influence of MPD to control perspective distortion is carried out via parameter $c$ and given probability $P$ to which MPD is applied during training.
@@
Objective 1- Robustness Evaluation of Existing models’ : Existing trained models need to be investigated for robustness against perspective distortion by evaluating them on the newly developed benchmark dataset ImageNet-PD.
@@
Objective 3- MPD’s effects on self-supervised learning: Investigation of the effect of MPD in learning robust self-supervised representations against perspective distortion.
@@ 3 Experiments & Results
Naming convention and experimental details: MPD defines the proposed transformation (shown in Fig. 1). In addition, we also have another variant, MPD IB, which replaces the black background with boundary foreground pixel values of the transformed image (shown in Fig. 14 in suppl. material). ImageNet-PD benchmark dataset has eight subsets (Fig. 3); four corresponding to four orientations (left, right, top, bottom) with black background (PD-L, PD-R, PD-T, PD-B). The other four subsets have the same orientations with interpolated backgrounds (PD-LI, PD-RI, PD-TI, PD-BI).
@@ 3.1 Existing models’ robustness evaluation
Figure 3: Perspectively distorted image examples from ImageNet-PD benchmark dataset. (a) Original image, (b) Left view (PD-L), (c) Right view (PD-R), (d) Top view (PD-T), (e) Bottom view (PD-B), (f) Left view with interpolation (PD-LI), (g) Right view with interpolation (PD-RI), (h) Top view with interpolation (PD-TI), (i) Bottom view with interpolation (PD-BI).
@@ 3.1.1 New benchmark dataset- ImageNet-PD:
To introduce perspective distortion by mimicking through the proposed transform MPD and its variant MPD-IB, we developed a new benchmarking dataset, ImageNet-PD (Fig. 3), derived from the ImageNet validation set. ImageNet-PD has eight subsets (Fig. 3), four corresponding to four orientations (left, right, top, bottom) with black background (PD-L, PD-R, PD-T, PD-B). The other four subsets are with the same orientations but with interpolated backgrounds (PD-LI, PD-RI, PD-TI, PD-BI).
Figure 4: Top1 and Top5 accuracies of standard architectures while tested on distorted images with and without background interpolation (BI). Interpolated images replace the black pixels in the distorted images with the boundary pixels; it makes images relatively more realistic than distorted images without interpolation (with surrounding black regions).
@@
In our evaluation across various deep learning architectures, from classical CNNs like ResNet, VGG, and AlexNet to modern structures like EfficientNet and Vision Transformers (ViT), we consistently observe a significant drop in model accuracy when faced with images exhibiting perspective distortions, refer Fig. 4 (also in Tables 6, 7 in suppl. material). AlexNet shows a drop from 56.52% to 37.25% in Top-1 and from 79.06% to 61.04% in Top-5; EfficientNet-b2 falls from 80.60% to 66.71% in Top-1 and from 95.31% to 86.30% in Top-5; and even Vision Transformer models like VIT-B-16 decline from 81.07% to 74.09% in Top-1 and from 95.32% to 90.98% in Top-5, as demonstrated by activation maps for a standard ResNet50 (Fig. 5, row 1). The uniform decrease in Top-1 and Top-5 accuracies across various models on ImageNet-PD versus ImageNet validation set shows perspective distortion as a real challenge in deep learning.

### v2

@@ Abstract
Perspective distortion (PD) causes unprecedented changes in shape, size, orientation, angles, and other spatial relationships of visual concepts in images. Precisely estimating camera intrinsic and extrinsic parameters is a challenging task that prevents synthesizing perspective distortion. Non-availability of dedicated training data poses a critical barrier to developing robust computer vision methods. Additionally, distortion correction methods make other computer vision tasks a multi-step approach and lack performance. In this work, we propose mitigating perspective distortion (MPD) by employing a fine-grained parameter control on a specific family of Möbius transform to model real-world distortion without estimating camera intrinsic and extrinsic parameters and without the need for actual distorted data. Also, we present a dedicated perspectively distorted benchmark dataset, ImageNet-PD, to benchmark the robustness of deep learning models against this new dataset. The proposed method outperforms existing benchmarks, ImageNet-E and ImageNet-X. Additionally, it significantly improves performance on ImageNet-PD while consistently performing on standard data distribution. Notably, our method shows improved performance on three PD-affected real-world applications—crowd counting, fisheye image recognition, and person re-identification—and one PD-affected challenging CV task: object detection. The source code, dataset, and models are available on the project webpage at https://prakashchhipa.github.io/projects/mpd.
@@
Perspective Distortion Self-supervised Learning Robust Representation Learning
@@ 1 Introduction
Perspective distortion (PD) in real-world imagery is omnipresent and poses challenges in developing computer vision applications. PD arises from various factors, including camera positioning, depth, intrinsic parameters such as focal length and lens distortion, and extrinsic parameters such as rotation and translation. These factors collectively influence the projection of 3D scenes onto 2D planes [42], altering semantic interpretation and local geometry. Precisely estimating these parameters for perspective distortion correction is challenging, presenting a critical barrier to developing robust computer vision (CV) methods.
@@
Previous studies in distortion correction predominantly concentrated on the correction of PD, with limited emphasis on enhancing the robustness of CV applications. Consequently, these correction methods transform CV tasks, such as image recognition and scene understanding, into a two-stage process: firstly, rectifying the distortion, and secondly, engaging in task-specific learning. GeoNet[29] employs CNN to predict distortion flow in images without prior distortion-type knowledge. PCL[57] uses perspective crop layers in CNNs to correct perspective distortions for 3D pose estimation. A cascaded deep structure network[46] corrects wide-angle portrait distortions without calibrated camera parameters. Another study[62] predicts per-pixel displacement for face portrait undistortion. PerspectiveNet[25] and ParamNet[25] predict perspective fields and derive camera calibration parameters, respectively. Methods for fisheye image distortion correction are also developed[54, 56]. Zolly[51] corrects perspective distortion by adjusting focal lengths in human mesh reconstruction tasks.
@@
Since the correction methods are fundamentally not focused on robust representation learning for CV tasks, they essentially follow an inefficient two-step process. One possible approach is to focus on developing task-specific methods where PD is noticeable, such as crowd counting[55, 22], object re-identification[2], autonomous vehicle[52], and recognizing texts in scene[39]. Further, [9] meta-learning based approach to adapt 3D human pose estimation to varying degrees of camera distortion. SPEC[26] tackled PD by estimating the camera parameters directly from in-the-wild images to improve the accuracy of 3D human pose predictions. These task-specific approaches fall short in providing generalizable and robust representation learning for PD mitigation, a crucial step for advancing CV research towards distortion robustness.
@@
The core reason behind the lack of generalizable and robust representation learning CV methods to mitigate PD is the absence of a real-world PD-affected training dataset. Both capturing real-world PD data and synthesizing PD-affected image data using parameter estimation are far from reality. Although existing benchmarks (ImageNet-X [23], ImageNet-E [28]) are partially affected by PD, a dedicated benchmark is needed. Therefore, there is a need to artificially synthesize PD in existing data without estimating camera intrinsic and extrinsic parameters. Since PD is non-linear in nature, the focus should be shifted from affine to potential non-linear transformation.
@@
One potential non-linear family of transformation is Möbius transform [1, 37] in complex theory. Due to its conformal nature (preserving angles during transformation), [63] utilized the Möbius transform for eight specific transformations, including various twists and spreads, as a data augmentation strategy to enhance model performance on standard data distributions. In other works, Möbius convolution[36], the authors design a Möbius-equivalent convolution operator in spherical CNN for shape classification and image segmentation tasks. OmniZoomer[5] integrates Möbius transformations into a neural network, enabling movement and variable zoom on omnidirectional images while addressing the challenges of blur and aliasing associated with such transformations. MöbiusGCN[3] addresses the limitation of existing graph convolution network GCNs in explicitly encoding transformations between joints in 3D human pose estimation. None of these works focus on perspective distortion.
@@
Due to the non-linear characteristics of Möbius transforming geometry in complex space while ensuring conformality, it is a potential approach to imitate perspective distortion when adapted to image data. Therefore, in the proposed mitigating perspective distortion (MPD) method, we mathematically model PD in the images by employing fine-grained parameter control on Möbius. It is governed by four parameters $a,b,c$ and $d$ (Eq. 1). MPD focuses on a specific family of transforms corresponding to the parameter $c$ to synthesize PD artificially. The core contribution of our method, MPD, is controlled parametrization for orientation (Fig. 1) and varying intensity (Fig. 2), which is capable of representing real-world perspective distortion.
@@
MPD transforms input image coordinates into complex vectors and performs a PD-specific family of Möbius transforms. Transformed complex vectors are then remapped to real-valued pixels, later discretized to obtain images mimicking perspective distortion while maintaining the integrity of visual concepts. Optionally, padding is applied to replace the black background with boundary pixel values to achieve another variant called integrated padding background variant (refer Fig. 15 in suppl. material). Notably, MPD achieves this without estimating intrinsic and extrinsic parameters, thus bypassing the need for actual distorted data.
@@
Incorporating PD in computer vision is beneficial for robustness since it can be used as an augmentation, resulting in improved performance on different computer vision tasks. To address the non-availability of a suitable benchmark, we have developed a dedicated perspectively distorted benchmark dataset, ImageNet-PD, derived from the ImageNet[11] validation set to evaluate the robustness of deep learning models against perspective distortion. We showcase a lack of robustness of standard deep learning models against perspective distortion by benchmarking them on ImageNet-PD.
@@
We employ ImageNet-PD in our proposed MPD method and extensively investigate it across supervised and self-supervised learning methods (SimCLR [7], DINO [6]). MPD demonstrates improved performance on existing benchmarks, ImagNet-E [28], and ImageNet-X [23]. MPD improves $10\%$ on ImageNet-PD while consistently maintaining the performance on the original ImageNet validation set. MPD shows performance improvements when adapted to multiple real-world CV applications. Specifically, MPD advances crowd counting with a novel method, MPD-AutoCrowd, achieving the highest performance (to the best of our knowledge) with a mean absolute error of 50.81 on ShanghaiTech-Part-A[60] and 96.80 on UCFCC50[21]. It demonstrates effective knowledge transfer in fish-eye image recognition, improving performance by $3\%$ on the VOC-360[14] dataset. MPD combined with a transformer-based self-supervised method, Clip-ReIdent[15], excels in person re-identification, achieving mean average precision of 97.02 and 98.30 with and without re-ranking, respectively, on the DeepSportradar dataset[48]. Also, MPD adapted for object detection with FasterRCNN [43] and FCOS [47] methods gain up to 3% performance improvement on COCO [30] dataset. The main contributions of this paper are as follows:
@@
    We develop a dedicated perspectively distorted dataset (ImageNet-PD) derived from the ImageNet[11] validation set to benchmark the robustness of computer vision models against PD to address the non-availability of a suitable benchmark.
@@ 2 Methodology
Conformality: Conformal transformations preserve angles between curves. The angle between two curves at a point is given by the argument of the complex ratio[19] of their derivatives. Let $f(z)$ and $g(z)$ represent curves in complex space. Their Möbius transforms are $\Phi(f(z))$, and $\Phi(g(z))$. If transformation preserves this ratio, it preserves angles. Eq. 3 shows that the angles are preserved, hence conformal [37]. Theoretical details and theorem are in suppl. material ( refer sec. 5 and 6).
@@ 2.1 Mitigating perspective distortion (MPD)
The proposed MPD method mathematically models perspective distortion in the image using a family of Möbius transform. First, it maps the discrete coordinates $(x,y)$ of the input image $I$ from pixel space to complex vector $z=x+iy$ in complex space. Second, it applies Möbius transform to $z$ and achieves corresponding non-linear transformed representation $z_{m}=x_{m}+iy_{m}$ by notably controlling real and imaginary parts of parameter $c$ in Eq. 1. Next, the transformed complex vector $z_{m}$ is mapped to real-valued pixel space $(x_{m},y_{m})$, which is discretized to get discrete pixel coordinates $(x_{d},y_{d})$. Obtained discrete pixel coordinates provide transformed image $I_{MPD}$ mimicking perspective distortion. Optionally, padding is applied to replace black background by boundary pixel to obtain integrated padding variant of transformed image. In this manner, MPD can synthesize perspective distortion artificially without estimating intrinsic and extrinsic parameters of perspective distortion. Noticeably, It eliminates the need for perspectively distorted real image data.
@@
The core part of MPD in modeling the PD lies in controlling the transformation through parameter $c$. Parameter $c$, being a complex number has real ($c_{real}$) and imaginary ($c_{imag}$) components; it is noteworthy that these components introduce real-world varieties of PD in the input image. $c_{real}$ perspectively distorts the in input image horizontally whereas $c_{imag}$ perspectively distorts the in input image vertically. The sign of $c_{real}$ dictates the direction ($+ve$ for left view and $-ve$ for right view) of distortion, whereas magnitude defines the intensity of applied distortion. Similarly, the sign of $c_{imag}$ dictates the direction ($+ve$ for top view and $-ve$ for bottom view) of distortion, whereas magnitude defines the intensity of applied distortion. The impact of signs of $c_{real}$ and imaginary $c_{imag}$ is depicted in Fig. 1. The synthesis of the increasingly distorted left view of the input image is demonstrated in Fig. 2 by controlling the intensity of $c_{real}$ component of MPD. Demonstrations of increasingly distorted right, top, bottom, and other views are provided in Fig. 17 in the suppl. material. The main contribution of our MPD method is its ability to control parameterization for varying intensity and orientation, effectively representing real-world perspective distortions. One of the ways to incorporate MPD is augmentation, which is investigated in this work. MPD is effortlessly employable to incorporate into the learning process to make networks robust against perspective distortion, materializing the proposed approach to shifting the focus from exclusion to inclusion of PD. The above-described method MPD is presented in Algorithm 1. MPD returned a transformed image of a ’cat’ as shown in step 11. Four regions are marked with colors (green, red, blue, and yellow) to observe the correspondence between input and transformed images. Although the current work does not focus on other parameters ($a$, $b$, and $d$), their impact on transformation is worth mentioning. The real component of parameters $a$ and $d$ controls scaling, and the imaginary components control rotations. The real component of parameter $b$ controls horizontal translations, and the imaginary component controls vertical translations. Their visualizations are provided in Fig. 18 in the suppl. material. The transformation parameters are set as: $a=1.0+0.0i$, $b=0.0+0.0i$, $d=1.0+0.0i$. The real and imaginary components of $c$ range from $0.2$ to $0.3$ to mimic perspective distortion. These values are uniformly applied in all experiments unless otherwise stated. The influence of MPD to control perspective distortion is carried out via parameter $c$ and given probability $P$ to which MPD is applied during training.
@@
  • Objective 1- Robustness evaluation of existing models: Investigation of existing trained models for robustness against perspective distortion by evaluating them on the PD-dedicated newly developed benchmark ImageNet-PD.
@@
  • Objective 3- MPD’s effects on self-supervised learning: Investigation of the effect of MPD in learning robust self-supervised representations against PD.
@@ 3 Experiments & Results
Naming convention and experimental details: MPD defines the proposed transformation (shown in Fig. 1). In addition, we also have another variant, MPD IB (integrated padding background), which replaces the black background with boundary foreground pixel values of the transformed image (shown in Fig. 16 in suppl. material). ImageNet-PD benchmark dataset has eight subsets (Fig. 3); four corresponding to four orientations (left, right, top, bottom) with black background (PD-L, PD-R, PD-T, PD-B). The other four subsets have the same orientations with integrated padding background, namely PD-LI (PD Left -Integrated padding background), PD-RI, PD-TI, and PD-BI.
@@ 3.1 Existing models’ robustness evaluation
Figure 3: Perspectively distorted image examples from ImageNet-PD benchmark dataset. (a) Original image, (b) Left view (PD-L), (c) Right view (PD-R), (d) Top view (PD-T), (e) Bottom view (PD-B), (f) Left view with integrated padding background (PD-LI), (g) Right view with integrated padding background (PD-RI), (h) Top view with integrated padding background (PD-TI), (i) Bottom view with integrated padding background (PD-BI).
@@ 3.1.1 New benchmark dataset- ImageNet-PD:
To introduce perspective distortion by mimicking through the proposed transform MPD and its variant MPD-IB, we developed a new benchmarking dataset, ImageNet-PD (Fig. 3), derived from the ImageNet validation set. ImageNet-PD has eight subsets (Fig. 3), four corresponding to four orientations (left, right, top, bottom) with black background (PD-L, PD-R, PD-T, PD-B). The other four subsets have the same orientations but integrated padding backgrounds using boundary pixels (PD-LI, PD-RI, PD-TI, PD-BI).
Figure 4: Top1 and Top5 accuracies of ImageNet trained models with standard architectures on ImageNet-PD subsets. Blue bars shows performance on original ImageNet validation set, green bars shows mean performance on ImageNet-PD subsets with black background, and orange bars shows mean performance on ImageNet-PD subsets with integrated padding background.
@@
In our evaluation across various deep learning architectures, from classical CNNs like ResNet, VGG, and AlexNet to modern structures like EfficientNet and Vision Transformers (ViT), we consistently observe a significant drop in model accuracy when faced with images exhibiting perspective distortions, refer Fig. 4 (also in Tables 7, 8 in suppl. material). AlexNet shows a drop from 56.52% to 37.25% in Top-1 and from 79.06% to 61.04% in Top-5; EfficientNet-b2 falls from 80.60% to 66.71% in Top-1 and from 95.31% to 86.30% in Top-5; and even Vision Transformer models like VIT-B-16 decline from 81.07% to 74.09% in Top-1 and from 95.32% to 90.98% in Top-5, as demonstrated by activation maps for a standard ResNet50 (Fig. 5, row 1). The uniform decrease in Top-1 and Top-5 accuracies across various models on ImageNet-PD vs. original ImageNet shows perspective distortion as a real challenge in computer vision.